3x2(3x to the second power) + 2x - 5 = 2.
the equation is 3x to the 2nd power not 3x2. Please help me solve.

Answer by Scriptor(36)

(3x)^2 + 2x - 5 = 2
9x^2 + 2x - 7 = 0
D = 4 - 4*9*(-7) = 4 + 252 = 256 = 16*16
x1 = (-2 + 16)/9 = 14/ 9
x2 = (-2 - 16) / 9 = -2